OpenStudy (anonymous):

Is there a function of two variables whose z=0 level curve consists of the circles x^2+y^2=4 and x^2+y^2=10? If so, what is an example? If not, why not?

OpenStudy (anonymous):

@Callisto

OpenStudy (anonymous):

@genius12

OpenStudy (anonymous):

\[f(x,y)=\left\{ \begin{array}{lr} 0, & (x^2+y^2==4) \vee (x^2+y^2==10) \\ 1, & (x^2+y^2\not=4) \wedge (x^2+y^2\not=10) \end{array} \right.\] Take a look at \(f(x)=x^4-4x^2+k,~~~0<k<4\).

OpenStudy (anonymous):

The first one is almost cheating, but it works. The second seems that , with some manipulation, and the introduction of a second independent variable \(y\), it could possible form two circles on the intersection of \(z=0\) and \(z=f(x,y)\).

OpenStudy (anonymous):

Yeah the first one definitely seems that way

OpenStudy (anonymous):

but i can see how that works though

OpenStudy (anonymous):

thank you